OpenStudy (anonymous):

Find a polynomial equation with real coecients that has −2 and i√2 as roots.

zepdrix (zepdrix):

Hey April :) If \(\rm x=-2\) is a root of some polynomial, adding 2 to each side of this equation gives us \(\rm x+2=0\). A polynomial which has -2 as a root will have \(\rm (x+2)\) as a factor! What we're doing is... we're using the `roots` to form `factors`. We'll then multiply these factors together to form our final `polynomial`.

zepdrix (zepdrix):

Recall that complex roots always come in `pairs`, conjugate pairs to be more precise. So if \(\rm x=i\sqrt2\) is a root, then \(\rm x=-i\sqrt2\) is also a root of the same polynomial.

zepdrix (zepdrix):

We'll do the same addition or subtraction trick from before, giving us this,\[\rm x-i \sqrt2=0,\qquad\qquad x+i \sqrt2=0\]So again, this gives us factors, \(\rm (x-i\sqrt2)\) and \(\rm (x+i\sqrt2)\), which will allow us to build up our polynomial.

zepdrix (zepdrix):

So to get your polynomial, multiply these factors together,\[\large\rm (x+2)(x-i \sqrt2)(x+i \sqrt2)\quad=\quad ?\]